Check your Understanding

Solve the following. Write your solutions on a separate sheet of paper.

1. A private organization distributed 5-kilo rice packs to a certain barangay placed under

special lockdown due to COVID 19. If the rice sells at P40 per kilogram, how much did the

organization spent if a total of 125 rice packs of rice was given?

2. An express delivery driver earns P500 daily for an eight-hour shift and an additional P35 for

every hour after that. Generate a function to solve his total earnings.

3. A computer shop charges P15 for the first hour and P10 for the succeeding hours. If Jacob

pays P45, how many hours did he rent the computer?

4. Ichan bought a car, he puts Php25 000 down payment and make monthly payment of Php12

000. If his car cost Php 457 000, how many years will it take him to pay it off?

5. Riza earns Php90 an hour in a toy factory with an additional of Php45 for her daily meal

allowance. On a Tuesday she earned a total of Php585, how many hours did she work for

that day?​


1
Expert's answer
2021-12-15T08:45:25-0500

1.

"125\cdot5\cdot40=P\ 25000"


2.

daily earning:

"y=500+35x"

where x is hours of overtime


3.

"45=15+3\cdot10"

Answer: "1+3=4" hours


4.

"25000+12\cdot12000x=457000"


"x=\frac{457000-25000}{144000}=3" years


5.

"90x+45=585"


"x=\frac{585-45}{90}=6" hours